Grade 12th passMechanics

The value of gravitational acceleration ‘g’ at a height ‘h’ above the earth’s surface is 4
g then
(R = radius of earth) THEN h=?

HI
8 Years agoGrade 12th pass

2 Answers

Phanindra
8 Years ago
We know that....g at ht. h , g(h) = g ÷ (1+h/R )^2.... Actually this is the formula...As one goes from bottom to top....Acceleration due to gravity decreases...But according to the question ... g(h) value is more than g...Can you please recheck the question...I will be glad to help you.... I am getting negative answer...
Arun
8 Years ago
Dear student
 
As we know
g(h) = g (1 – 2h / R)
now
4g = g (1 – 2h / R)
4 – 1 = – 2h / R
h =  – 3R /2
 
Regards
Arun